LOFAR Newsletters August 2016

Distributed on 29 August 2016

-------------------------------------

 

Announcements:

----------------

- The Cycle 7 proposal call can be found here.

- Starting from 31 August 2016, the LOFAR Status Meetings will have a monthly cadence

- Because of necessary maintenance activities, between 1-16 October 2016, LOFAR data stored at SURFsara can not be accessed at all. This outage applies to data links that include "grid.sara.nl" in the path. To understand whether a given Cycle/DDT project is stored at SARA, please check the 'Details' section within your project folder in MoM and navigate to 'Resources'.

 

Array status:

----------------------

- 38 stations operational in the Netherlands: 24 core stations and 14 remote stations.

- 12 international stations operational: DE601, DE602, DE603, DE604, DE605, FR606, SE607, UK608, DE609, PL610, PL611, PL612 

- A new LOFAR station (IE613) will be built on the grounds of Birr Castle, located centrally in Ireland. Expected completion: first half 2017.

- Oscillating tiles were detected during BF observations at a few stations and were deactivated.

- The maintenance is in progress and will continue till the weather conditions will allow.

- The overview of non-operational antenna elements for LBA and HBA is available here.

-  Station calibration:

o   Efforts toward the calibration of the LBA SPARSE mode are still going on and imply software adjustments. Data for LBA INNER have been collected: manual data flagging is needed for almost all stations. New calibration tables for HBA LOW have been produced and COBALT delays for the core stations are about to be produced as well.

 

 

Observing System Status:

------------------------------------ 

- The observing system experienced instability during the period July - August 2016. Problems were associated with the processing and LTA (ingest) systems and COBALT nodes. These caused delays in the exploitation of the Cycle 6 observing programmes and in digesting data out of CEP2.

- The new CEP4 cluster is under commissioning. The imaging observing mode is almost fully commissioned - a few final tests are in progress. Commissioning of the Beamformed and TBB modes is still in progress. CEP4 commissioning has priority over Cycle observing. Due to this, a few production observations have been shifted to the next Cycle.

 

 

Software Development Status (J. Annyas):

------------------------------------------

- CEP4 commissioning is still being supported.

- The new LCUs have arrived. The LCUs + OS and software versions are being tested and, after the results are clear, a rollout plan for this year will be made.

- The first version of Survey Tier I ingest was delivered.

 

 

CITT Status (E. Orru' & T. J. Dijkema)

--------------------------------------------------------------------

- There will be soon a release of the prefactor pipeline (direction independent)

- The factor pipeline (direction dependent) reached an advanced level of automatization and flexibility. Discussions are ongoing for having it released.

- DPPP: the new algorithm to fit for TEC has been tested during the busy week. Some issues have been identified and are now under investigation.

- A pilot test showed that the new image domain gridder runs properly on CEP4. Development is on going to tight the new solver to wsclean.

- Baseline dependent averaging has been implemented and tested on wsclean. This showed an improvement in processing speed of about one order of magnitude.

 

 

MSSS Status (J. Broderick & G. Heald)

--------------------------------------------------------------------

- There has been good progress recently towards the next version of the MSSS-HBA catalog. A field-based flux scale correction method has yielded very encouraging results, with much more accurate in-band spectral indices across the full survey area. Further quality control checks are being carried out.   

- A campaign to reprocess the HBA data at higher angular resolution will commence in the coming months. In preparation, benchmarking is being performed on CEP3 and other external clusters.

 

 

Observing Programmes

--------------------------------------------------------------------

- Cycle 6 observing programme: 45% complete. The observing schedule can be fund here.

- Cycle 5 observing programme: 92% complete. The rest is being observed with second priority during Cycle 6.

 

Documentation

--------------------------------------------------------------------

- The list of LOFAR papers has been updated - it is available here.

- The LOFAR Visibility Calculator has been released and can be fund here.

 

 

CEP news:

----------------

-       CEP4
  • The CEP4 cluster is being commissioned.

-       CEP2:

  • Disk space situation is often critical, especially in case of instability of the processing and ingest systems and after week-ends. 
  • During the past two months, a few locus nodes experienced issues and were temporarily taken out of production.

-       CEP3:

  • Cluster info and schedule available here.
  • Monitoring of node usage is active. Statistics will be used to evaluate requests for extensions.

 

 

Calendar next LOFAR activities:

-------------------------------------------------------------------

Note: the following events are now marked on an online calendar that is available here.

- LOFAR Data Processing School: 5-9 September 2016

- Next LSM's : 31/08, 28/09, 26/10

(all presentations given at the LSM and video recordings are available here).

- Next Stop days: 30/08, 04/10, 01/11

- Next software roll outs: 07/11

- Next LOFAR bulletin: October 2016

Design: Kuenst.    Development: Dripl.    © 2020 ASTRON